# Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES)

### v2

@@ III-A Robot Architecture
The Furhat robot was employed in the study, featuring a neutral-looking face that underwent user validation before interactions. The robot’s face engine incorporated smiles and eyebrow raises during conversations to enhance naturalness and provide non-verbal feedback to users without context analysis. To further refine the interactions, the robot incorporated subtle behaviors like blinking, eye shifts, and brief gaze aversion while speaking, based on silences in user input.
@@
GPT-3.5 (text-davinci-003, OpenAI) was used for dialogue generation, as it was the most capable LLM at the time (March 2023). Prompting was used to give an empathetic persona to the robot, guiding it to ask open questions, listen actively with follow-up questions, and reflect on situations.
@@
Initially, English was chosen as the communication language for speech recognition (Google Cloud Speech-to-Text), dialogue generation, and synthesis (Amazon Polly) due to the more extensive training data available for LLMs. However, an initial study with Swedish-speaking older adults showed the need for communicating in their native language [12]. Hence, Swedish was used in the follow-up study investigated in this paper. A USB microphone array (Seeed Studio) was used in both studies to obtain clear audio for speech recognition.
@@
While the conversation with the robot was autonomous, a wizard interface was used to start the interaction with the user by entering the participant ID. The initial and final222Robot response to start the interaction: “Hello! I am Furhat, the personalized companion robot. What is your name?” Robot response to end the interaction: “I would love to talk more another time, but for the sake of time, I need to say goodbye. Thank you for talking with me. Take care!” robot responses were pre-scripted to ensure that the interaction started and ended the same way for all participants. The rest of the interaction was fully autonomous based on the user’s responses and the responses generated by the LLM. The wizard interface was also used to end the interaction if necessary, i.e., if the participant wants to end the conversation early or an error occurs in the system that requires a restart to continue the conversation where it is left off. A 7-minute timer was set (checked automatically after each user response), after which the robot would say its pre-scripted response to ensure a fair comparison between users.
III-B Data
To understand older adults’ perceived benefits and challenges in interaction with a conversational companion robot, and correspondingly develop a robot that meets their expectations, we conducted preliminary interviews with Swedish-speaking older adults aged 65 and over, in which they talked with a robot autonomously and individually in English. The study had 6 (3 men, 3 women) Swedish-speaking healthy older adults, between 66 to 86 years old (, ). However, due to it being the very first evaluation of an LLM on a social robot with older adults, the interactions had a lot of failures that are analyzed thoroughly in [12]. Interaction failures can lead to lower likeability and satisfaction [44] and cause negative tone and emotion in user responses [45]. Thus, to prevent biasing the enjoyment scale solely towards negative experiences, but also have an understanding of user reactions to frequent technical failures in current architectures, we chose the most successful interaction (as defined by the least number of failures and the longest length of interaction) from this study to be included as a basis of alignment for developing the enjoyment scale, as described in Sections IV-B and IV-C. The other criterion was that the interaction contained ‘highs and lows’, that is, the participant reacted positively (e.g., smile, laugh), neutrally, and negatively (e.g., frown, getting impatient) in the video to enable the annotators to understand the spectrum of responses. The corresponding subject (denoted as S0) was an 83-year-old male without prior experience of robots. The participant lived with their partner in their own home. The interaction lasted 13.5 min (53 turns). The video was recorded from a side angle, facing both the participant and the robot. The participant gave informed consent for recording, analysis of data, and anonymized (blurred and without a name) image and video sharing for publications.
@@ III-B Data
Following the preliminary interviews, technical improvements were made for turn-taking, the robot’s persona, and architecture to overcome the interaction failures. Based on the feedback from the interviews, the interaction language was changed to Swedish. Subsequently, four participatory design workshops were conducted with 28 older adults having an autonomous open-domain conversation with the robot individually for approximately 7 minutes. Prior to the robot interactions, the robot’s capabilities were demonstrated through a researcher having a conversation with the robot (2 min), and focus group discussions were made using design scenarios of everyday activities to understand their expectations of companion robots. The researcher(s) were present in the room (to interfere if necessary) during the individual robot interactions. Following the interactions, the participants completed a 68-question Likert scale (1 to 5) questionnaire based on HRI and open-domain dialogue literature. Based on several studies described in Section II, user satisfaction, fun, and interestingness of the conversation were evaluated and categorized under the user enjoyment construct. To account for discomfort in the conversation [46], the strangeness of the conversation was also evaluated, similar to [47]333“Did you feel something strange in that dialogue with the robot?” was used in [47]. We adapted it slightly for the Likert scale., by reverse-coding it in analysis:
@@
All interactions with the robot were video-recorded by an external camera facing both the participant and the robot at a side angle, as well as through the robot camera to record the participant’s face. All participants gave informed consent for recording, analysis of data, and anonymized (blurred and without a name) image and video sharing for publications.
@@
Two of the subjects were selected for annotator alignment: a 69-year-old woman (denoted as S1) and a 75-year-old man (denoted as S2). S1 and S2 did not have any prior interaction with a robot. S1 and S2 lived with their partners in their own homes. The selection basis was to find interactions containing a range of ‘highs and lows’, as in the previous study, aiming to complement S0 with interactions with fewer failures. S1’s interaction lasted 7.5 min (27 turns) and S2’s interaction lasted 7.3 min (27 turns).
After the initial alignment of annotators with 3 videos to create HRI CUES (Section IV), the remaining participant interactions (except one due to lack of a side-video) were used for HRI CUES evaluation (Section V). The resulting data consisted of 25 participants’ (12 men, 13 women) interactions, with a mean age of 74.6 (). 20 participants had no prior interaction with a robot, and only one had previously talked with a robot. Interaction duration was min () with 12 to 29 turns. Each turn lasted 5 to 61 seconds (, ). The total duration of the videos was 174 min, corresponding to 590 turns.
We provide the dataset for HRI CUES that includes anonymized transcripts, annotation scores, and self-reported user perceptions online444HRI CUES Dataset: https://paperswithcode.com/dataset/hri-cues-dataset. Videos of the interactions are available upon request, contingent upon a signed agreement to maintain data confidentiality in accordance with General Data Protection Regulation.
@@ IV Assessing Enjoyment from Conversations
This work addresses the lack of user enjoyment analysis of conversation from an external perspective in HRI. We start from an existing enjoyment scale in human-human relations to develop HRI CUES, by complementing it with annotations of older adults’ interactions with a conversational companion robot. This section describes not only the scale proposed in this work, but also a final complete methodology to evaluate enjoyment from videos of conversations with robots. It also provides annotation guidelines and details the practices taken for establishing inter-rater reliability in annotations [48].
@@ IV-B Familiarizing with Data
As a starting point to familiarize annotators with the data and develop a user enjoyment scale for conversational HRI, the annotators were given a slightly adapted version of the enjoyment scale by Reimnitz and Rauer5551 (very low): no evidence of pleasure. Pair never has fun or enjoys the interaction, although there may be joint interaction. There is no mutual enjoyment of positive affect or negative interaction. 3 (neutral anchor): there is occasional positivity that is not strong or frequently displayed and may be displayed by only one partner towards the other. Pair is doing OK together but without real joy or enthusiasm for their shared interactions. 7 (very high): the pair is very satisfied with the interaction and activity. The couple shows mutual enjoyment in their interaction marked with shared exuberance and/or delight. There is consistent visual regard coupled with affective sharing. [16] on human-human conversations of married couples, in which ‘user’ was used instead of ‘couple’, and references that relate to couples (‘mutual enjoyment’, ‘affective sharing’, and ‘exuberance’) were removed. Instead of a 7-point Likert scale, which may be difficult to align given only the lowest, natural anchor, and highest enjoyment values, a 5-point scale was used:
@@
Due to the subjectivity of user enjoyment, it was necessary to establish common grounds on the levels of user enjoyment prior to the annotators analyzing all the robot interactions individually [48]. Thus, three exemplar videos (S0, S1, and S2) as explained in Section III-B were chosen that contain a range of negative and positive responses from the user and a variety of technical failures.
@@
Conversational turns (exchanges) were chosen as the basis of annotations, because they were mostly similar in duration for participants, as well as for paving the way for understanding user enjoyment via autonomous systems to adapt and improve the interaction continuously. As such, the segments to be annotated were created automatically based on the turns (Robot-Participant pairs) in manually-corrected (for timing and content) transcripts. All videos started with the robot’s first phrase (greeting of the user). A turn ends (and a new turn starts) when the participant stops speaking, as that holds the potential to be detected by an automatic system for evaluating the turn that could be used to generate a new response.
Annotators were guided to apply the rating scale on a per-turn basis, assessing both the robot’s response and the participant’s subsequent input within each turn. Furthermore, they were tasked with delivering an overarching assessment of enjoyment encompassing the entire interaction, referred to as overall enjoyment. Within this context, annotators were encouraged to provide an in-depth rationale for their ratings, adopting an open-ended approach to offer comprehensive insights. They were asked to elaborate on the aspects and multimodal cues they considered in shaping their evaluations, along with the methodology they employed. Additionally, they were asked to provide whether any challenges or difficulties were encountered while evaluating overall user enjoyment for the interaction and its details. The ratings for each turn were to be recorded within the ELAN file (Fig. 2), while a separate document was designated for annotators to record their overall interaction rating and provide open-ended responses.
@@ IV-C Annotator Alignment
Annotators were requested to systematically review the three videos, examining each conversation turn individually. They were prompted to assess various factors, including rating, cues, and aspects, while also considering the detectability of user enjoyment in each turn. Annotators were further instructed to identify contrasting and supporting arguments for their ratings, as the reasons behind the divergence between the annotators can be just as, if not more, valuable than the concordance between the annotators [62]. Upon completion of the turn analysis, annotators were guided to discuss their overall user enjoyment rating, identify which aspects, multimodal cues, and conversation turns contributed most significantly to their conclusions, and consider the relative importance (weights) assigned to these modalities in their assessments. Annotators were also encouraged to evaluate whether specific aspects or cues were observable from both the frontal view (the robot’s camera that directly captures participants) and the side view (the external camera that captures both entities) and reflect on how analyzing these different perspectives may have influenced their assessments. Annotators were advised to keep their ratings, both within ELAN and the accompanying document that justifies their scores, readily accessible on their screens as a reference point during the discussions.
Self-reported user perceptions were given to the annotators at this stage, as reported by questionnaire ratings in terms of the level of user satisfaction, fun, interestingness of the conversation, and strangeness of talking to the robot, as described in Section III-B. These metrics correlate highly with each other (Cronbach’s ), with lower when any of the metrics are excluded. The annotators were instructed to view the ratings after the overall user enjoyment in the interaction had been discussed for the corresponding participant. Based on the user perceptions, how these results correlate with their findings and the reasons behind discrepancies were discussed.
@@ IV-D Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES)
The discussions between the annotators (A1, A2, and A3) lasted four hours, during which they carefully went through three example videos (S0-S2) that they had previously annotated and discussed the cases one by one. The annotators viewed every turn in the interaction several times and exchanged the reasons behind their rating in terms of multimodal cues, until they all aligned on the rating for each turn. Simultaneously, they created a list of signs of enjoyment and dis-enjoyment that they had used during their enjoyment evaluation. Towards the end of the session, they settled on a 5-item scale based on the initial provided scale, ranking from very low enjoyment to very high enjoyment. The final user enjoyment scale, namely the Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES), is:
@@
Two exchanges per rating of the scale from the alignment interactions are provided as illustrative examples in the video provided in the footnote666HRI CUES exemplary exchanges: https://youtu.be/VmKvGM0pyec.
To rate an exchange higher on the user enjoyment scale (4 or 5), the annotators looked for different signs of enjoyment, which, for example, were smirking, movement, flow of conversation (the topic is moving forward), no strain or discomfort, asking questions [to the robot], smooth turn-taking, dynamic tonality (and dynamic phrasing of sentences), being playful, sharing personal experiences [to the robot], sharing an understanding (common ground) [with the robot], and anthropomorphizing [the robot].
To rate an exchange lower on the scale (1 or 2), the annotators looked for signs of dis-enjoyment, which, for example, were low energy, sighing, tiredness, long breaths, restless movements (i.e., adaptors, such as moving in the chair from side to side or changing arm position), flat tonality, silence, awkward and negative facial expressions, flaring nostrils, disengagement cues (e.g., turning away from the robot, or responding in a way that disrupts the conversation flow, such as “That is true”), and topic closure (e.g., “Let’s talk about something else”). In addition, robot behaviors that disrupted the interaction flow, such as repeated questions, were considered to be strong causes of dis-enjoyment.
@@
In cases where the exchange has cues from multiple scale levels, the annotators determined the dominant level in that interaction. This could be done by observing the intensity of the cues, the significance of the cues, or the interaction trajectory. On the other hand, if an annotator observed strong cues from two moderately or highly distinct levels (as opposed to subsequent levels), they would annotate using a level between those. For instance, as evident in this exchange777Exchange with cues from multiple levels: https://youtu.be/2HA-_5B9JHs, when there is discomfort at the beginning (1), but the user continues to politely keep up the interaction (3), the exchange would be annotated as a 2, the mid-point between the levels.
@@
When annotating the videos, the annotators assumed that in the future, robots would be able to judge the level of user enjoyment in real-time while having a conversation. Therefore, the videos were annotated segment by segment (turn by turn), with each segment being watched only once, similar to [40].
IV-E Annotation using HRI CUES

### v3

@@ III-A Robot Architecture
The Furhat robot was employed in the studies, featuring a neutral-looking face that underwent user validation before interactions. The robot’s face engine incorporated smiles and eyebrow raises during conversations to enhance naturalness and provide non-verbal feedback to users without context analysis. To further refine the interactions, the robot incorporated subtle behaviors like blinking, eye shifts, and brief gaze aversion while speaking, based on silences in user input.
@@
GPT-3.5 (OpenAI, initial study: text-davinci-002, second study: text-davinci-003) was used for dialogue generation, as it was the most capable LLM at the time (September 2022 / March 2023). Prompting was used to give a persona to the robot. Initially, this focused on getting to know more about the conversation partner, however, this created superficial conversations that were not interesting for older adults. That is why, the persona was changed for the latter study to be empathetic, guiding it to ask open questions, listen actively with follow-up questions, and reflect on situations.
@@
Initially, English was chosen as the communication language for speech recognition (Google Cloud Speech-to-Text), dialogue generation, and synthesis (Amazon Polly) due to the more extensive training data available for LLMs. However, the initial study with Swedish-speaking older adults showed the need for communicating in their native language. Hence, Swedish was used in the follow-up study.
A USB microphone array (Seeed Studio) was used in both studies to obtain clear audio for speech recognition. A silence-based threshold was used for turn-taking, where the robot would take the turn (generate a response and synthesize it) after a fixed period of silence following the end of the user’s speech, without listening to the user while taking the turn. In the initial study, the participants were interrupted frequently. Thus, this threshold was increased for the second study and a red-light signal was added underneath the robot to signal the robot was taking the turn, which substantially decreased the interruptions.
@@
While the conversation with the robot was autonomous, a wizard interface was used to start the interaction with the user by entering the participant ID. The initial and final222Robot response to start the interaction: “Hello! I am Furhat, the personalized companion robot. What is your name?” Robot response to end the interaction: “I would love to talk more another time, but for the sake of time, I need to say goodbye. Thank you for talking with me. Take care!” robot responses were pre-scripted to ensure that the interaction started and ended the same way for all participants. The rest of the interaction was fully autonomous based on the user’s responses and the responses generated by the LLM. The wizard interface was also used to end the interaction if necessary, i.e., if the participant wants to end the conversation early or an error occurs in the system that requires a restart to continue the conversation where it is left off. In the initial study, the participant ended the interaction whenever they wanted (by saying “Goodbye”). In contrast, in the second study, a 7-minute timer was set (checked automatically after each user response), after which the robot would say its pre-scripted response to ensure a fair comparison between users.
III-B Study Details
Ethical approval was granted by the Swedish Ethical Review Authority (reference number 2022-09-21) for the following two studies in [10]. All participants provided informed consent for data recording, analysis, and the use of anonymized (blurred and nameless) images and videos in publications.
III-B1 Preliminary Study
Preliminary interviews were conducted with Swedish-speaking older adults aged 65 and over, in which they talked with a robot autonomously and individually in English, for 4 to 13 minutes.
The study had 6 (3 men, 3 women) Swedish-speaking healthy older adults, between 66 to 86 years old (, ).
III-B2 Second Study
@@
Following the preliminary interviews, technical improvements were made for architecture, persona prompt, language, and turn-taking to overcome the interaction failures, as summarized in Section III-A. Subsequently, a second study was conducted with 28 older adults having an autonomous open-domain conversation with the robot individually for approximately 7 minutes. Prior to the robot interactions, the robot’s capabilities were demonstrated through a researcher having a conversation with the robot (2 minutes), and focus group discussions were made using design scenarios of everyday activities to understand their expectations of companion robots. The researcher(s) were present in the room (to interfere if necessary) during the individual robot interactions. Following the interactions, the participants completed a 68-question Likert scale (1 to 5) questionnaire in [10], based on HRI ([35, 47, 48, 49, 50, 51, 52, 53]) and open-domain dialogue ([54, 55, 56]), ranging from constructs on user perceptions (e.g., enjoyment, ease of use, usefulness, anxiety towards robots) to capabilities of the agent (e.g., turn-taking, consistency, fluency of dialogue). For the purposes of this work, only the user enjoyment construct was used, which focuses on user satisfaction, fun, and interestingness of the conversation [35, 47]), also accounting for discomfort in the conversation [57] with the strangeness of the conversation, similar to [48]333“Did you feel something strange in that dialogue with the robot?” was used in [48]., by reverse-coding it in analysis:
@@
The Cronbach’s alpha (for the evaluation data) was , showing high correlation of items with each other in the construct, with the removal of any item decreasing .
All interactions with the robot were video-recorded by an external camera facing both the participant and the robot at a side angle, as well as through the robot camera to record the participant’s face.
@@
III-C Alignment Data
Three videos were chosen for the initial alignment of annotators to create HRI CUES (Sections IV-B to IV-E).
Due to the preliminary study being the very first evaluation of an LLM on a social robot with older adults, the interactions had a lot of failures, which were analyzed thoroughly in [10]. Interaction failures can lead to lower likeability and satisfaction [58] and cause negative tone and emotion in user responses [59]. Thus, to prevent biasing the enjoyment scale solely towards negative experiences, but also have an understanding of user reactions to frequent technical failures in current architectures, we chose the most successful interaction (as defined by the least number of failures and the longest length of interaction) from this study to be included as a basis of alignment for developing the enjoyment scale, as described in Sections IV-B and IV-C. The other criterion was that the interaction contained ‘highs and lows’, that is, the participant reacted positively (e.g., smile, laugh), neutrally, and negatively (e.g., frown, getting impatient) in the video to enable the annotators to understand the spectrum of responses. The corresponding subject (denoted as S0) was an 83-year-old male without prior experience of robots. The interaction lasted 13.5 minutes (53 turns). The video was recorded from a side angle, facing both the participant and the robot.
In addition, two of the subjects from the second study were selected for annotator alignment: a 69-year-old woman (denoted as S1) and a 75-year-old man (denoted as S2). S1 and S2 did not have any prior interaction with a robot. The selection basis was to find interactions containing a range of ‘highs and lows’, as in the previous study, aiming to complement S0 with interactions with fewer failures. S1’s interaction lasted 7.5 minutes (27 turns) and S2’s interaction lasted 7.3 minutes (27 turns).
Conversational exchanges (i.e., a robot turn followed by a participant turn - termed here as a turn for brevity) were chosen as the basis of annotations, because they were mostly similar in duration for participants, as well as for paving the way for understanding user enjoyment via autonomous systems to adapt and improve the interaction continuously. As such, the segments to be annotated were created automatically based on the turns in manually-corrected (for timing and content) transcripts. All videos started with the robot’s first phrase (greeting of the user). An exchange ends (and a new exchange starts) when the participant stops speaking, as that holds the potential to be detected by an automatic system for evaluating the exchange that could be used to generate a new response.
III-D Evaluation Data
The remaining participant interactions from the second study (except one due to lack of a side-video) were used for HRI CUES evaluation (Sections IV-F and V). The resulting data consisted of 25 participants’ (12 men, 13 women) interactions, with a mean age of 74.6 (). 20 participants had no prior interaction with a robot, and only one had previously talked with a robot. Interaction duration was minutes () with 12 to 29 turns. Each exchange lasted 5 to 61 seconds (, ). The total duration of the videos was 174 minutes, corresponding to 590 turns. The videos were segmented using conversational turns, as explained above.
We provide the evaluation dataset for HRI CUES that includes anonymized transcripts, annotation scores, and self-reported user perceptions online [60]444HRI CUES Dataset [60]: https://doi.org/10.5281/zenodo.12588810. Videos of the interactions are available upon request, contingent upon a signed agreement to maintain data confidentiality in accordance with General Data Protection Regulation.
@@ IV Assessing Enjoyment from Conversations
This work addresses the lack of user enjoyment analysis of conversation from an external perspective in HRI. Thus, we started from an existing enjoyment scale in human-human relations to develop HRI CUES, by complementing it with annotations of older adults’ interactions with a conversational companion robot. This section describes not only the scale proposed in this work, but also a complete methodology to evaluate enjoyment from videos of conversations with robots. It also provides annotation guidelines and details the practices taken for establishing inter-rater reliability in annotations [61].
@@ IV-B Familiarizing with Data
As a starting point to familiarize annotators with the data and develop a user enjoyment scale for conversational HRI, the annotators were given a slightly adapted version of the enjoyment scale by Reimnitz and Rauer5551 (very low): no evidence of pleasure. Pair never has fun or enjoys the interaction, although there may be joint interaction. There is no mutual enjoyment of positive affect or negative interaction. 3 (neutral anchor): there is occasional positivity that is not strong or frequently displayed and may be displayed by only one partner towards the other. Pair is doing OK together but without real joy or enthusiasm for their shared interactions. 7 (very high): the pair is very satisfied with the interaction and activity. The couple shows mutual enjoyment in their interaction marked with shared exuberance and/or delight. There is consistent visual regard coupled with affective sharing. [18] on human-human conversations of married couples, in which ‘user’ was used instead of ‘couple’, and references that relate to couples (‘mutual enjoyment’, ‘affective sharing’, and ‘exuberance’) were removed. Instead of a 7-point Likert scale, which may be difficult to align given only the lowest, neutral anchor, and highest enjoyment values, a 5-point scale was used:
@@
Due to the subjectivity of user enjoyment, it was necessary to establish common grounds on the levels of user enjoyment prior to the annotators analyzing all the robot interactions individually [61]. Thus, three exemplar videos (S0, S1, and S2) as explained in Section III-C were chosen that contain a range of negative and positive responses from the user and a variety of technical failures.
@@
Annotators were guided to apply the rating scale on a per-exchange basis, assessing both the robot’s response and the participant’s subsequent input within their turn. Furthermore, similar to [13], they were tasked with delivering an overarching assessment of enjoyment encompassing the entire interaction, referred to as overall enjoyment. Within this context, annotators were encouraged to provide an in-depth rationale for their ratings, adopting an open-ended approach to offer comprehensive insights. They were asked to elaborate on the aspects and multimodal cues they considered in shaping their evaluations, along with the methodology they employed. Additionally, they were asked to provide whether any challenges or difficulties were encountered while evaluating overall user enjoyment for the interaction and its details. The ratings for each exchange were to be recorded within the ELAN file (Fig. 2), while a separate document was designated for annotators to record their overall interaction rating and provide open-ended responses.
@@ IV-C Annotator Alignment
Annotators were requested to systematically review the three videos, examining each conversation exchange individually. They were prompted to assess various factors, including rating, cues, and aspects, while also considering the detectability of user enjoyment in each exchange. Annotators were further instructed to identify contrasting and supporting arguments for their ratings, as the reasons behind the divergence between the annotators can be just as, if not more, valuable than the concordance between the annotators [75]. Upon completion of the turn-by-turn analysis, annotators were guided to discuss their overall user enjoyment rating, identify which aspects, multimodal cues, and conversation turns contributed most significantly to their conclusions, and consider the relative importance (weights) assigned to these modalities in their assessments. Annotators were also encouraged to evaluate whether specific aspects or cues were observable from both the frontal view (the robot’s camera that directly captures participants) and the side view (the external camera that captures both entities) and reflect on how analyzing these different perspectives may have influenced their assessments. Annotators were advised to keep their ratings, both within ELAN and the accompanying document that justifies their scores, readily accessible on their screens as a reference point during the discussions.
Self-reported user perceptions were given to the annotators at this stage, as reported by questionnaire ratings in terms of the level of user satisfaction, fun, interestingness of the conversation, and strangeness of talking to the robot, as described in Section III-B2. These metrics demonstrate high reliability with each other (Cronbach’s ), with lower when any of the metrics are excluded. The annotators were instructed to view the ratings after the overall user enjoyment in the interaction had been discussed for the corresponding participant. Based on the user perceptions, how these results correlate with their findings and the reasons behind discrepancies were discussed.
@@
| Signs of disenjoyment | Signs of enjoyment |
| Low energy/Tiredness | Smirking |
| Sighing | Engaged movement (e.g. leaning in) |
| Long breaths | No strain or discomfort asking questions [to the robot] |
| Restless movements | Smooth turn-taking |
| Flat tonality | Dynamic tonality |
| Silence | Sharing personal experiences [to the robot] |
| Awkward and negative facial expressions | Sharing an understanding |
| Flaring nostrils | Anthropomorphizing [the robot] |
| Disengagement cues | Asking questions to the robot |
| Topic closure | Being playful |
| Repeated questions | Flow of conversation |
@@ IV-D Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES)
The discussions between the annotators (A1, A2, and A3) lasted four hours, during which they carefully went through three example videos (S0-S2) that they had previously annotated and discussed the cases one by one. The annotators viewed every exchange in the interaction several times and exchanged the reasons behind their rating in terms of multimodal cues, until they all aligned on the rating for each exchange. Simultaneously, they created a list of signs of enjoyment and dis-enjoyment that they had used during their enjoyment evaluation, which can be seen in Table I. Towards the end of the session, they settled on a 5-item scale based on the initial provided scale, ranking from very low enjoyment to very high enjoyment. The final user enjoyment scale, namely the Human-Robot Interaction Conversational User Enjoyment Scale (HRI CUES), is:
@@
To rate an exchange higher on the user enjoyment scale (4 or 5), the annotators looked for different signs of enjoyment (see Table I), which, for example, were flow of conversation (i.e., the topic is moving forward), dynamic tonality, and dynamic phrasing of sentences, as well as sharing an understanding, i.e., having common ground with the robot.
To rate an exchange lower on the scale (1 or 2), the annotators looked for signs of dis-enjoyment (see Table I), which, for example, were restless movements i.e., adaptors, such as moving in the chair from side to side or changing arm position. Disengagement cues e.g., turning away from the robot, or responding in a way that disrupts the conversation flow, such as “That is true”), and topic closure (e.g., “Let’s talk about something else”) is also an example of disenjoyment. In addition, robot behaviors that disrupted the interaction flow, such as repeated questions, were considered to be strong causes of dis-enjoyment.
@@
In cases where the exchange has cues from multiple scale levels, the annotators determined the dominant level in that interaction. This could be done by observing the intensity of the cues, the significance of the cues, or the interaction trajectory. On the other hand, if an annotator observed strong cues from two moderately or highly distinct levels (as opposed to subsequent levels), they would annotate using a level between those. For instance, as evident in this exchange666Exchange with cues from multiple levels: https://youtu.be/2HA-_5B9JHs, when there is discomfort at the beginning (1), but the user continues to politely keep up the interaction (3), the exchange would be annotated as a 2, the mid-point between the levels.
@@
When annotating the videos, the annotators assumed that in the future, robots would be able to judge the level of user enjoyment in real-time while having a conversation. Therefore, the videos were annotated segment by segment (turn-by-turn), with each segment being watched only once, similar to [12].
| Turn | Exchange | Reason | Score |
| 11 |
ROBOT: Oh, I just wasn’t sure what you meant. Are you asking why I didn’t wanna go with you?
PARTICIPANT: No, I asked where you want to go. Do you have a favorite place? |
Misunderstanding, gaze aversion, sighing, low energy | 1 |
| 8 |
ROBOT: I am sorry, I didn’t understand you.
PARTICIPANT: Okay. I will have a walk in nature reserve in.. |
Interruption, rushed speech, gaze aversion, boredom | 2 |
| 18 |
ROBOT: That sounds great! I love nature too.
PARTICIPANT: What kind of nature do you like? |
Polite continuation, awkward expression | 3 |
| 23 |
ROBOT: Mount Everest is about 8848 meters high.
PARTICIPANT: Oh, good. Thanks. I didn’t know it that exactly. |
Smooth interaction, nodding, high energy | 4 |
| 22 |
ROBOT: I can try.
PARTICIPANT: How high is Mount Everest? |
Quick response, smiling, leaning forward, immersion | 5 |
IV-E Examples for the Scale
During alignment, annotators discussed each exchange in the three videos in detail to distinguish between different levels of the scale. Two exchanges per rating of the scale from the alignment interactions are provided as illustrative examples in the video provided in the footnote777HRI CUES exemplary exchanges: https://youtu.be/VmKvGM0pyec. Table II presents five of these example exchanges for one participant (S0), along with the corresponding scores and reasoning.
In Turn 11, the participant sounded disappointed that the robot misunderstood the question and repeats their previous inquiry. He looks at the experimenter three times (gaze aversion) and sighs before posing their second question for clarification. The participant displays low energy throughout. These behaviors indicate signs of frustration. In Turn 8, the robot interrupts the participant at the beginning of the turn. In response, the participant raises their head in a slow nod, signaling annoyance, and looks at the experimenter (gaze aversion). He then repeats the sentence he was previously attempting to say in a rushed manner. The participant inhales to finish the sentence, which are signs associated with boredom. In Turn 18, the robot interrupts the participant at the beginning of the exchange, causing the participant to stop speaking and make an awkward expression. The robot’s phrasing (“That sounds X. I love Y too”) is repeated for the fifth time in the conversation, with “I love nature too” repeated for the third time. The participant nods, smiles, and asks a follow-up question in a polite manner to move the conversation forward. In Turn 23, the interaction proceeds smoothly with no interruptions or communication failures. The participant appears interested in the conversation, as the robot correctly interprets and responds to their question. The participant nods in affirmation and responds with high energy. In Turn 22, the robot interrupts the participant mid-speech. Despite the interruption, the participant exhibits rapid turn-taking in response, demonstrating immersion in the conversation. He smiles, swing sideways in their chair, and lean forward while asking the next question, which are signs that reflect excitement.
While these examples are intended to illustrate what may constitute a given level on the scale according to the signs of enjoyment relevant to the use case in this work (Table I), the interpretation can depend on the conversation context and history (e.g., prior robot failures may cause user frustration), vary between participants (e.g., extroverts vs. introverts, older adult vs. young adult), rely on the dominant signals in the exchange (e.g., mixed signals with differing intensities), and differ across application domains (e.g., games may elicit deeper immersion than casual interactions with a companion robot). As highlighted in the previous section, it is important for annotators to establish the ‘baseline’ behavior from which an individual may deviate during the interaction, while considering what has happened within the conversation history. Section VI-C provides guidelines for adapting the scale to other domains through expert review of the relevant signals of enjoyment for each level of the scale.
IV-F Annotation using HRI CUES